OpenStudy (s-mlipinsky):

Find funtion g(x) given that g(g(x)) = 6x-8

OpenStudy (sooobored):

oh, what an interesting problem im going to assume g(x) will have the form g(x)=ax^b +c hence g(g(x))= a (ax^b +c) ^b +c hmmm, no clue

OpenStudy (yellowfellow007):

if in a g(x) ax+b format, im almost positive a has to equal root 6.

OpenStudy (518nad):

what is g(g(g(x)))

OpenStudy (518nad):

what is g(g(g(g(x))))

OpenStudy (s-mlipinsky):

What would the -8 be?

OpenStudy (518nad):

6(6x-8)-8=g(g(g(g(x)))) 36x-42-8=36x-50=gggg(x)

OpenStudy (yellowfellow007):

yo i solved it. its \[\sqrt{6}x- 2.31918\]

OpenStudy (yellowfellow007):

give that a try

OpenStudy (518nad):

g(x)=ax+b form right so we need root6

OpenStudy (yellowfellow007):

what i figured is -8=-root6x-x. solved for that and got it

OpenStudy (518nad):

yeah

OpenStudy (yellowfellow007):

i just tried it again and it seemed to work

OpenStudy (yellowfellow007):

:D good job guys

OpenStudy (518nad):